Circular No. 7414

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

MARKARIAN 421 D. J. Fegan and S. Dunlea, University College, Dublin; H. Badran,

M. Catanese, S. Fegan, D. Horan, and T. C. Weekes, Smithsonian Astrophysical Observatory; J. Buckley, Washington University; J. Finley, Purdue University; F. Krennrich, Iowa State University; and J. Rose, University of Leeds, on behalf of the Whipple Observatory Gamma-Ray Collaboration, report significantly increased TeV-γ-ray emission from Markarian 421 during Apr. 27–May 1: “All observations were taken above an elevation of 50^{o} with the 490-pixel Imaging Atmospheric Cerenkov Telescope, at an operating threshold of ∼ 275 GeV. Preliminary analysis yields detected rates, expressed in units of the Crab-nebula TeV rate, as follows: Apr. 26, 0.32 ± 0.18; 27, 1.47 ± 0.07; 28, 1.24 ± 0.07; 29, 0.98 ± 0.07; 30, 2.90 ± 0.08; May 1, 1.22 ± 0.08. Observations taken on Apr. 30 are particularly interesting, exhibiting an almost uniform increase in event rate over a 2-hr period, from a baseline value of < 0.4 Crab at the start to a rate of 4.7 Crab at the maximum of the flare. Thereafter, the flare decayed slowly over a further 2-hr period, reaching a rate of 2.4 Crab when tracking observations ceased.”

SUPERNOVA 2000cc IN CGCG 140-014 G. Aldering and A. Conley, Lawrence Berkeley National Laboratory,

write: “An optical spectrum (range 330–900 nm, with S/N ∼ 11/nm) of SN 2000cc (cf. IAUC 7413), taken with the Cerro Tololo 4-m telescope (+ RC spectrograph) on May 2.4 UT, exhibits a featureless spectrum, consistent with that of a blackbody of ∼ 8500 K. This object may prove to be a supernova at a very early epoch, and we encourage further monitoring.”

SUPERNOVA 2000bk IN NGC 4520 Further to their item on IAUC 7408, Jha et al. report that their spec-

trum of SN 2000bk does not show interstellar Na I D absorption from the host galaxy; the Na I D feature reported there was misidentified.

SUPERNOVA 2000cb IN IC 1158 K. Krisciunas, University of Washington; and N. C. Hastings and R.

McMillan, Apache Point Observatory (APO), report preliminary photometry (uncertainities ± 0.02 mag) of SN 2000cb, using the APO 3.5-m telescope (+ software aperture of diameter 4′′.5): May 1.41 UT, V = 17.96, B − V = +0.88, V − R = +0.41, V − I = 0.60.

2000 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green